### v1

@@ 1 Main
Figure 2: Proposed fingerprinting strategy and associated task distance measure: binned Kullback-Leibler Divergence (bKLD). Top: To compute a task fingerprint, $n=10,000$ images are sampled (I) and passed through a pretrained backbone (we use an ImageNet [2] ResNet34 [3]) to extract $m=512$ features per image (II). The resulting $m\times n$ features are binned into $b=100$ bins along the features axis (III), resulting in m (normalized) histograms representing the fingerprint. Bottom: To compare two fingerprints, a weighted sum of the Kullback-Leibler Divergence (KLD) across all per-feature histograms is computed. A softmax operation is applied to source task histograms to avoid empty bins.
The concept of Transfer Learning dates back to the 1970s [4] and involves any knowledge transfer from a source task to a target task [5], with its most prominent application being pretraining on large-scale datasets [6, 7]. While utilizing off-the-shelf pretrained models is a common practice to speed up model convergence [6], the vast number of available architectures, pretraining schemes, and datasets, makes selecting the most suitable option a time-consuming process. This premise also holds for the nascent field of foundation models [8]. It has been shown that simply relying on benchmarks for pretrained models, such as, ImageNet [2], does not translate well to the medical domain [6]. This fueled research of Task Transferability Estimation (TTE), which aims to quantify the potential of knowledge transfer between tasks [9]. While the research problem has been known for decades [10] and substantial progress has been made, challenges with respect to data privacy considerations, robustness in realistic heterogeneous data settings, and the avoidance of negative transfer remain [11, 7, 12]. Existing transferability estimation methods, which attempt to assess model suitability for a target task, often have only been applied in unrealistically homogeneous settings [13, 14, 15], lack large-scale validation (12 tasks in [16], 7 tasks in [17]), or are incompatible with data privacy requirements [18]. Others are not suited for scalability, either because of computational complexity or the assumption that tasks share underlying images [19, 20, 21, 22].
@@
Figure 4: Task fingerprinting benefits training pipeline configuration and beats manual knowledge transfer. Top: Fraction of n=43 validation tasks that improve Balanced Accuracy (BA) through knowledge transfer (”gain” [20]) in four scenarios. Bottom: Average delta in Area Under the Receiver Operator Characteristic (AUROC) across n=43 validation tasks. X-axis shows the number of shots, translating to the best of top k suggestions of our framework. Error bars correspond to standard deviation over three repetitions of all model trainings. Our proposed binned Kullback-Leibler Divergence (bKLD) fingerprint (here: the small variant) improves training for up to 90% of validation tasks.
@@
Figure 5: binned Kullback-Leibler Divergence (bKLD) outperforms previously proposed methods for knowledge transfer. Uncertainty-aware ranking of our proposed bKLD methods for task fingerprinting versus VisualDNA (VDNA) [17], Predict To Learn (P2L) [25], Fisher Embedding Distance (FED) [26, 14], Fréchet Inception Distance (FID) [27, 28] and manual task selection. Bottom: Columns represent four meta metrics to compare distance measures, whilst rows correspond to four knowledge transfer scenarios. We average across the top three suggestions by each method (weightedtau evaluates the full candidate ranking). Blob size shows frequency of rank across 1000 bootstrap samples from 258 setups (2 base metrics, 43 validation tasks, 3 repetitions). X marks the mean rank and whiskers the standard deviation. Plot is inspired by [29] and follows the “aggregate then rank” assessment method. Top: Summary of the 16 subplots below. The marker position refers to the mean rank over individual bootstrapped mean rankings (X marks) and whiskers indicate standard deviation.
@@ 2.1 Task fingerprinting enables knowledge transfer without sharing sensible data
For a fraction of 67% of validation tasks, the resulting models exhibited an improvement by switching our baseline model architecture to the prime candidate proposed by the knowledge cloud. Replacing the augmentation strategy benefited 58% of tasks’ first shot, i.e., only leveraging the augmentation strategy derived from the best matching source task in the knowledge cloud. Using shared image data of the best-aligned source task for knowledge transfer through pretraining and co-training resulted in an improvement for 41% and 57% of validation tasks. Increasing the computational budget during downstream model training and considering multiple knowledge source candidates (”multi-shot”) led to up to 90% of validation tasks being improved, as shown in Fig. 4 (top). Notably, this held true when transferring model architecture (90%) and augmentation policy (83%) solely via sharing the fingerprint as a compressed identifier without any other data samples leaving the institution of the data owner. Assuming that the task data was available for pre- or co-training, we observed a similar rate of task improvement for the co-training scenario (84%), compared to our single task baseline. Considerably fewer tasks benefited from knowledge transfer via pretraining (75%), while the fluctuations in our repeated experiments due to randomness were more severe.
@@ 2.2 Task fingerprinting outperforms manual task selection
Reviewing the literature to identify previously successful training recipes has proven to be the predominant strategy chosen by participants of biomedical challenges [1]. This strategy, which we refer to as ”manual” task matching, solely relies on the similarity of the semantic task description, is limited to the actual training information published on a similar task, and requires manual extraction from the publication or repository. Simulating such semantic task matching allowed us to compare our fully data-driven task fingerprinting with this prevailing strategy. The results are shown in Fig. 4 (bottom), which displays the average performance increase per task compared to our single task baseline. The relative performance increase by task fingerprinting was 12%, 57%, 15%, and 2% larger than manual task matching for the investigated knowledge transfer scenarios of model architecture, pretraining, augmentation policy, and co-training, respectively, when averaging across all tasks, repetitions, and the multi-shot range from 1 to 5.
@@ 2.3 Our proposed task distance measure outperforms previously suggested ones
To compare our fingerprinting technique to previously proposed task similarity measures, we instantiated our framework with previously proposed task transferability measures including VisualDNA (VDNA) [17], Predict To Learn (P2L) [25], Fisher Embedding Distance (FED) [26, 14] and Fréchet Inception Distance (FID) [27, 28]. Note that the underlying task representations differ for each of these, with FID and VDNA being comparable in size to our bKLD-based task fingerprints, while P2L is significantly smaller and FED is larger. We compare them to three variants (hyperparameter configurations) of our proposed distance measure bKLD, which obtain the top three positions in the overall ranking (see Fig. 5 top). The bottom of Fig. 5 shows a detailed analysis of the rankings for all mentioned methods with respect to meta metris and knowledge transfer scenarios. Within all four knowledge transfer scenarios, for at least three out of four meta metrics, a single variant of bKLD performed best. Additional evaluation aggregation schemes can be found in the supplementary information (Supplementary Table LABEL:tab:win_rates and Supplementary Fig. LABEL:fig:app_comparison).
Figure 6: binned Kullback-Leibler Divergence (bKLD) is robust with respect to dataset size. Percentile of best 3-shot source task knowledge transfer according to Balanced Accuracy (BA) averaged over n=43 validation tasks in four scenarios applying three proposed bKLD variants. X-axis shows the number of samples used from a task to generate its fingerprint in log scale. Error bars indicate standard deviation over 10 resamplings.
Comparing transferability measures has been shown to be a very sensitive experiment, where small variations in the setup may lead to conflicting conclusions on superiority [30]. Therefore, the setup stability score [30], has been proposed to indicate how stable the ranking of task similarity algorithms is with respect to varying a single component of the setup, e.g. choosing a different performance metric. A value of 0 indicates no correlation between the rankings and 1 denotes perfectly identical rankings. For our experiments, the stability scores were 0.61 for the meta metrics, 0.32 for the performance metrics, 0.12 for the random seed, 0.02 for the target tasks, and only 0.01 for the knowledge transfer scenario.
@@ 2.4 Our proposed task distance measure is robust with respect to task size
As fingerprint generation is a non-deterministic process and we envision our methodology to be especially useful for small tasks, we varied the number of samples used to compute a fingerprint. While all previous experiments (including all alternative task similarity measures) evaluated 10,000 samples from a task, we tested significantly reduced sizes of up to only 10 samples. Figure 6 shows our method’s stability in retrieving sufficiently beneficial task candidates in a best of three shots scenario, where random task selections would be expected to achieve a value of 0.75. Based on fingerprints generated from only 10 samples bKLD(large,unweighted) and bKLD(small,target) show robust task matching performance. Of note, knowledge transfer for pretraining data has the best relative task matching of the four scenarios investigated, meaning that out of the available sources typically ones around the 88th percentile are chosen (Fig. 6). This stands in contrast to the absolute improvement to the validation tasks as shown in Fig. 4, where applying pretraining knowledge transfer benefited the least across the scenarios investigated.
@@ 3 Discussion
    Task fingerprinting allows for secure knowledge transfer: Our findings demonstrate the potential of task fingerprinting as a method for knowledge sharing, particularly in the data-sensitive domain of medical imaging. Our approach improved downstream task performance for a majority of tasks within a reasonable computational budget, while outperforming the prevailing and time-consuming practice of manual knowledge compilation.
@@
    The best fingerprinting strategy varies substantially across scenarios: The reason we did not identify a one-size-fits-all solution is that the optimal source task rankings are inconsistent, hence no single task distance measure might suit all transfer scenarios. The weighting mechanism of bKLD solves this dilemma and shifts attention to features that are relevant for a given scenario.
@@
    Task distance measures benefit from feature binning: Our newly introduced task distance measure bKLD outperformed existing measures in all evaluated transfer scenarios. While the Predict To Learn (P2L) [25] approach computes the Kullback-Leibler Divergence (KLD) on averaged features, the binning strategy in bKLD allows a much more granular comparison of image feature distributions. The introduced weighting mechanism further provides an option for prioritization of dominant features of the source or target task.
@@
While our method is straightforward to use, it comes with one important hyperparameter; the weighting mechanism enables bKLD to adapt to various scenarios of knowledge transfer. Setting the weights of bKLD according to dominant source task features was especially beneficial in the model architecture scenario (see Fig. 5 bottom), though this variant of bKLD was the least robust with respect to image sampling (see Fig. 6) and the scenario obtained the lowest percentile score (see Fig. 6). Furthermore, manual task selection suited this scenario better than all others, presumably indicating that granular feature distributions have a subordinate role in the fitting of model architectures. For this specific scenario, also known as source-free model transferability estimation, many solutions have been proposed, but the problem remains unsolved [12]. For the pretraining scenario, bKLD(small,target) proposed source tasks ranked highest over all scenarios (see Fig. 6). Interestingly, in absolute terms, the performance improvement was low compared to the other training scenarios (see Fig. 4), i.e., for pretraining the bKLD method appeared to perform best, but the improvement is still small compared to other scenarios. We interpret this as a lack of tasks that are suited well for pretraining in our experimental data pool, as pretraining tasks ideally are large in size, and comprise sufficient variety in sample distribution to cover the downstream task [31, 32]. The consistently poor performance of bKLD(large,source) in this scenario opposed to the other weighting strategies (see Fig. 5 bottom) further reveals the importance of feature distribution similarity especially with respect to dominant features of the target task. In the co-training scenario, where the unweighted bKLD variant performed best, it appears intuitive that focusing primarily on one task’s dominant features does not yield better transfer estimation. Though the bKLD(small,target) and bKLD(large,unweighted) variants’ meta metrics were very close to each other in most cases (see Fig. 5, 6, Supplementary Fig. LABEL:fig:app_comparison).
Closely related to task fingerprinting is the field of Automated Machine Learning (AutoML), which aims to fully automate the configuration of ML algorithms, specifically steps such as data preprocessing, feature engineering, and hyperparameter optimization [33]. It shares the goal of AI democratization by resolving tricky steps in ML and empowers non-experts to generate AI systems. However, AutoML often raises computational expenses, especially in approaches such as Neural Architecture Search (NAS) [34, 35], and, in contrast to our approach, lacks a collaborative mechanism. Similarly, frameworks such as nnU-net [36] have shown promise in offering self-configuring solutions for biomedical image segmentation tasks. Although training adjustment in nnU-net is way faster and substantially less resource-intensive than traditional AutoML, it is limited to the static rules designed by the authors (i.e., no new knowledge is gained automatically over time), focuses on a specific task type (only segmentation) and excludes crucial aspects of the training pipeline (e.g., architecture template and data augmentation are fixed). Federated learning [37], on the other hand, offers a path for collaborative learning while preserving data privacy, but imposes additional burdens related to synchronization, trust establishment, and preparation [38]. Task fingerprinting combines the efficiency of nnU-net with the flexibility of AutoML, while leveraging collaborative efforts such as Federated Learning. Similarly to the latter, a general limitation of task fingerprinting is that, in a public setting, the system is vulnerable to malicious actors submitting false data, requiring the development of preventive measures. In addition, the success of the overall concept relies on the willingness of researchers to share knowledge within the community. Perhaps most importantly and in contrast to all three mentioned approaches, however, our method has the potential to facilitate Lifelong Learning [39] by continuously and automatically growing knowledge, as more data and tasks are incorporated over time. This capability could lead to long-term scalability of AI which would reduce energy consumption and carbon footprint of systems [40].
It is important to consider our findings in the context of previous research on task distance measures. In line with our preliminary work [26] the FED task selector ranks high for the model architecture and pretraining scenarios and some meta metrics. However, it falls short on the other meta metrics and knowledge transfer scenarios. Prior studies have also highlighted the inconsistency of transferability estimations in general [30] and specifically in the medical domain [41]. This aligns well with our observations regarding the low stability scores of existing distance measures across target tasks and the transfer scenario and emphasizes the importance of finding a flexible approach that can adapt to multiple transfer scenarios. In fact, Standley et al. [42], also observed that task similarity measures might be suited for one kind of knowledge transfer, e.g., pretraining, but inappropriate for another, e.g., co-training. Our aggregation over multiple metrics and meta metrics was intended to counteract the sensitivity of evaluations to specific metrics [43] and increase the assessment robustness as well as allow for the consideration of multiple knowledge transfer scenarios. We observed that weightedtau allows for the most stable separation of task selectors, as opposed to regret, percentile, and improvement (Fig. 5 bottom, Supplementary Fig. LABEL:fig:app_comparison). However, it is hard to interpret and circumvents a resource limitation-based cutoff.
The emergence of foundation models [8] has amplified the importance of AI democratization, as the requisite expertise, computational resources, and data volumes have grown substantially. Such foundation models might bridge the gap between different modalities in medical imaging through their large-scale pretraining [44]. However, they are accompanied by two fundamental problems: first, navigating the diverse landscape of available models to select the most suitable one, and second, devising an effective adaptation strategy for specific downstream tasks. While the integration of multiple modalities into fingerprints is left for future research, we note that these challenges are well-suited to be answered by task fingerprinting.

### v2

@@ 1 Main
Figure 2: Proposed fingerprinting strategy and associated task distance measure: binned Kullback-Leibler Divergence (bKLD). To compute a task fingerprint, $n=10,000$ images are sampled (I) and passed through a pretrained backbone (we use an ImageNet [2] ResNet34 [3]) to extract $m=512$ features per image (II). The resulting $m\times n$ features are binned into $b=100$ bins along the features axis (III), resulting in m (normalized) histograms representing the fingerprint. To compare two fingerprints, a weighted sum of the Kullback-Leibler Divergence (KLD) across all per-feature histograms is computed. A softmax operation is applied to source task histograms to avoid empty bins.
The concept of Transfer Learning dates back to the 1970s [4] and involves any knowledge transfer from a source task to a target task [5], with its most prominent application being pretraining on large-scale datasets [6, 7]. While utilizing off-the-shelf pretrained models is a common practice to speed up model convergence [6], the vast number of available architectures, pretraining schemes, and datasets, makes selecting the most suitable option a time-consuming process. This premise also holds for the nascent field of foundation models [8]. It has been shown that simply relying on benchmarks for pretrained models, such as, ImageNet [2], does not translate well to the medical domain [6]. This fueled research of Task Transferability Estimation (TTE), which aims to quantify the potential of knowledge transfer between tasks [9]. While the research problem has been known for decades [10] and substantial progress has been made, challenges with respect to data privacy considerations, robustness in realistic heterogeneous data settings, and the avoidance of negative transfer remain [11, 7, 12]. Existing transferability estimation methods, which attempt to assess model suitability for a target task, often have only been applied in unrealistically homogeneous settings [13, 14, 15, 16], lack large-scale validation (12 tasks in [17], nine tasks in [18], eight tasks in [19], seven tasks in [20], five tasks in [21]), or are incompatible with data privacy requirements [22]. Others are not suited for scalability due to computational complexity [23, 24, 25, 19, 26, 27] or the assumption that tasks share underlying images [28].
@@
Figure 4: Task fingerprinting benefits training pipeline configuration and beats manual knowledge transfer. (a) : Fraction of n=43 validation tasks that improve Balanced Accuracy (BA) through knowledge transfer (”gain” [28]) in four scenarios. (b) : Average delta in Area Under the Receiver Operator Characteristic (AUROC) across n=43 validation tasks. X-axis shows the number of shots, translating to the best of top k suggestions of our framework. Error bars correspond to standard deviation over three repetitions of all model trainings. Our proposed binned Kullback-Leibler Divergence (bKLD) fingerprint (here: the small variant) improves training for up to 90% of validation tasks.
@@
Figure 5: binned Kullback-Leibler Divergence (bKLD) outperforms previously proposed methods for knowledge transfer. Uncertainty-aware ranking of our proposed bKLD methods for task fingerprinting versus manual task selection, VisualDNA (VDNA) [20], Fisher Embedding Distance (FED) [31, 15], Fréchet Inception Distance (FID) [32, 33], Predict To Learn (P2L) [34], and Sinkhorn Divergence [35]. (a): Columns represent four meta metrics to compare distance measures, whilst rows correspond to four knowledge transfer scenarios. We average across the top three suggestions by each method (weightedtau evaluates the full candidate ranking). Blob size shows frequency of rank across 1000 bootstrap samples from 258 setups (2 base metrics, 43 validation tasks, 3 repetitions). X marks the mean rank and whiskers the standard deviation. Plot is inspired by [36] and follows the “aggregate then rank” assessment method. (b): Summary of the 16 subplots below. The marker position refers to the mean rank over individual bootstrapped mean rankings (X marks) and whiskers indicate standard deviation.
@@ 2.1 Task fingerprinting enables knowledge transfer without sharing sensible data
For a fraction of 67% of validation tasks, the resulting models exhibited an improvement by switching our baseline model architecture to the prime candidate proposed by the knowledge cloud. Replacing the augmentation strategy benefited 58% of tasks’ first shot, i.e., only leveraging the augmentation strategy derived from the best matching source task in the knowledge cloud. Using shared image data of the best-aligned source task for knowledge transfer through pretraining and co-training resulted in an improvement for 41% and 57% of validation tasks. Increasing the computational budget during downstream model training and considering multiple knowledge source candidates (”multi-shot”) led to up to 90% of validation tasks being improved, as shown in Fig. 4 (a ). Notably, this held true when transferring model architecture (90%) and augmentation policy (83%) solely via sharing the fingerprint as a compressed identifier without any other data samples leaving the institution of the data owner. Assuming that the task data was available for pre- or co-training, we observed a similar rate of task improvement for the co-training scenario (84%), compared to our single task baseline. Considerably fewer tasks benefited from knowledge transfer via pretraining (75%), while the fluctuations in our repeated experiments due to randomness were more severe.
@@ 2.2 Task fingerprinting outperforms manual task selection
Reviewing the literature to identify previously successful training recipes has proven to be the predominant strategy chosen by participants of biomedical challenges [1]. This strategy, which we refer to as ”manual” task matching, solely relies on the similarity of the semantic task description, is limited to the actual training information published on a similar task, and requires manual extraction from the publication or repository. Simulating such semantic task matching allowed us to compare our fully data-driven task fingerprinting with this prevailing strategy. The results are shown in Fig. 4 (b), which displays the average performance increase per task compared to our single task baseline. The relative performance increase by task fingerprinting was 12%, 57%, 15%, and 2% larger than manual task matching for the investigated knowledge transfer scenarios of model architecture, pretraining, augmentation policy, and co-training, respectively, when averaging across all tasks, repetitions, and the multi-shot range from 1 to 5.
@@ 2.3 Our proposed task distance measure outperforms previously suggested ones
To compare our fingerprinting technique to previously proposed task similarity measures, we instantiated our framework with previously proposed task transferability measures including VisualDNA (VDNA) [20], Predict To Learn (P2L) [34], Fisher Embedding Distance (FED) [31, 15], Sinkhorn Divergence [35], and Fréchet Inception Distance (FID) [32, 33]. Note that the underlying task representations differ for each of these, with FID and VDNA being comparable in size to our bKLD-based task fingerprints, while P2L is significantly smaller and FED as well as Sinkhorn being larger. We compare them to two variants (hyperparameter configurations) of our proposed distance measure bKLD, which obtain the top two positions in the overall ranking (see Fig. 5 b ). Fig. 5 (a) shows a detailed analysis of the rankings for all mentioned methods with respect to meta metrics and knowledge transfer scenarios. Additional evaluation aggregation schemes can be found in the supplementary information (Supplementary Table LABEL:tab:win_rates and Supplementary Fig. LABEL:fig:app_comparison).
Figure 6: binned Kullback-Leibler Divergence (bKLD) is robust with respect to dataset size. Percentile of best 3-shot source task knowledge transfer according to Balanced Accuracy (BA) averaged over n=43 validation tasks in four scenarios applying two proposed bKLD variants. X-axis shows the number of samples used from a task to generate its fingerprint in log scale. Error bars indicate standard deviation over 10 resamplings.
Comparing transferability measures has been shown to be a very sensitive experiment, where small variations in the setup may lead to conflicting conclusions on superiority [37]. Therefore, the setup stability score [37], has been proposed to indicate how stable the ranking of task similarity algorithms is with respect to varying a single component of the setup, e.g. choosing a different performance metric. A value of 0 indicates no correlation between the rankings and 1 denotes perfectly identical rankings. For our experiments, the stability scores were 0.59 for the meta metrics, 0.31 for the performance metrics, 0.1 for the random seed, 0.02 for the target tasks, and only 0.01 for the knowledge transfer scenario.
@@ 2.4 Our proposed task distance measure is robust with respect to task size
As fingerprint generation is a non-deterministic process and we envision our methodology to be especially useful for small tasks, we varied the number of samples used to compute a fingerprint. While all previous experiments (including all alternative task similarity measures) evaluated 10,000 samples from a task, we tested significantly reduced sizes of up to only 10 samples. Figure 6 shows our method’s stability in retrieving sufficiently beneficial task candidates in a best of three shots scenario, where random task selections would be expected to achieve a value of 0.75. Based on fingerprints generated from only 10 samples both bKLD variants show robust task matching performance. Of note, knowledge transfer for pretraining data has the best relative task matching of the four scenarios investigated, meaning that out of the available sources typically ones around the 88th percentile are chosen (Fig. 6). This stands in contrast to the absolute improvement to the validation tasks as shown in Fig. 4, where applying pretraining knowledge transfer benefited the least across the scenarios investigated.
Figure 7: binned Kullback-Leibler Divergence (bKLD) generalizes to 3D, new task types and further backbone models. Uncertainty-aware ranking for 2D classification and 3D medical image segmentation tasks demonstrating strong performance of our proposed bKLD methods for task fingerprinting compared to manual task selection, VisualDNA (VDNA) [20], Fréchet Inception Distance (FID) [32, 33], Predict To Learn (P2L) [34], and Sinkhorn Divergence [35]. (a): Extension of Fig. 5 to three additional feature extraction backbones. Summary of 48 subplots (16 from a CLIP [38], DINO [39], and MAE [40] backbone respectively). The marker position refers to the mean rank over individual bootstrapped mean rankings (X marks) and whiskers indicate standard deviation. (b): Extension of Fig. 5 to 3D medical imaging using data from the Medical Segmentation Decathlon [41]. Columns represent four meta metrics to compare distance measures. We average across the top three suggestions by each method (weightedtau evaluates the full candidate ranking). Blob size shows frequency of rank across 1000 bootstrap samples from 34 setups (2 base metrics, 17 validation tasks). X marks the mean rank and whiskers the standard deviation. Plot is inspired by [36] and follows the “aggregate then rank” assessment method.
2.5 Our proposed task distance measure generalizes to unseen task type and feature extractors
To further test the generalization capabilities of bKLD, we investigated two variations of our experiments. First, we applied task transferability estimation on the outcomes of the Medical Segmentation Decathlon [41]. In this international competition 19 participating teams submitted segmentation models for 17 segmentation tasks on 10 datasets. Data were obtained from the two most common 3D medical imaging modalities, namely 3D magnetic resonance imaging (MRI) and computed tomography (CT). Segmentation tasks were highly diverse, ranging from the segmentation of large objects (e.g., liver) to the segmentation of tubular structures (e.g., vessels). Note that the task type (segmentation versus classification), the modalities (2D versus tomographic 3D) and the target structures (e.g., vessels) all represent a severe shift compared to our previous experiments [42]. The outcomes are shown in Fig. 7 (b) and are consistent with our previous findings. Concretely, the small variant of bKLD performs best in three out of four meta metrics. Second, we replaced the feature extractor used in our previous experiments with three more recent models based on CLIP [38], DINO [39], and MAE [40]. These experiments represent shifts in the distribution of the extracted features and consequently the resulting task fingerprints. Fig. 7 (a) shows the aggregated pendant to Fig. 5 (b), i.e., the ranking of fingerprinting methods across meta metrics, transfer scenarios, and feature extractors, and is consistent with our main experiments.
@@ 3 Discussion
    Task fingerprinting allows for secure knowledge transfer. Our findings demonstrate the potential of task fingerprinting as a method for knowledge sharing, particularly in the data-sensitive domain of medical imaging. Our approach improved downstream task performance for a majority of tasks within a reasonable computational budget, while outperforming the prevailing and time-consuming practice of manual knowledge compilation.
@@
    The best fingerprinting strategy varies substantially across scenarios. The reason we did not identify a one-size-fits-all solution (see Fig. 5 a) is that the optimal source task rankings are inconsistent (see Supplementary Table LABEL:tab:scenario_similarity), hence no single task distance measure might suit all transfer scenarios. Despite this, bKLD demonstrates strong generalization capabilities.
@@
    Task distance measures benefit from feature binning. Our newly introduced task distance measure bKLD outperformed existing measures . While the Predict To Learn (P2L) [34] approach computes the Kullback-Leibler Divergence (KLD) on averaged features, the binning strategy in bKLD allows a much more granular comparison of image feature distributions. The introduced weighting mechanism further provides an option for prioritization of dominant features of the target task.
While our method is straightforward to use, it comes with some important hyperparameters. The number of bins allows to control the tradeoff between representation granularity on the one hand and privacy and efficiency on the other hand. The weighting mechanism allows to make use of the same representation in multiple ways, thus enables bKLD to adapt to various scenarios of knowledge transfer. Importantly, the comparability of fingerprints requires an identical feature extraction model and at least a similar preprocessing method. Otherwise the $j$-th feature dimension of model $A$ is unlikely to be directly related to the $j$-th feature dimension of model $B$, and the total number of feature dimensions may differ. Although our experiments (Fig. 7) demonstrate that bKLD generalizes well beyond a single feature extractor, major considerations regarding the maintainability of knowledge in the knowledge cloud remain. In theory, it would be feasible to support multiple fingerprinting pipelines (comprising preprocessing and feature extraction) in parallel for images with widely varying natures (e.g., histopathology and 3D imaging) and occasionally update these pipelines to support multiple extraction models over time. This approach would provide sufficient flexibility to introduce more powerful feature extractors while avoiding complete erasure of existing knowledge with the introduction of a single new extraction pipeline.
Compared with previous methods, bKLD performed poor in the model architecture scenario (see Fig. 5 a), which also obtained the lowest percentile score (see Fig. 6). Furthermore, manual task selection suited this scenario better than all others, presumably indicating that granular feature distributions have a subordinate role in the fitting of model architectures. For this specific scenario, also known as source-free model transferability estimation, many solutions have been proposed, but the problem remains unsolved [12]. On a wider scope our generalization experiment on the Medical Segmentation Decathlon shows that the transfer of a complete model pipeline may be better predicted (Fig. 7 b). For the pretraining scenario, bKLD(small) proposed source tasks ranked highest over all scenarios (see Fig. 6). Interestingly, in absolute terms, the performance improvement was low compared to the other training scenarios (see Fig. 4), i.e., for pretraining the bKLD method appeared to perform best, but the improvement is still small compared to other scenarios. We interpret this as a lack of tasks that are suited well for pretraining in our experimental data pool, as pretraining tasks ideally are large in size, and comprise sufficient variety in sample distribution to cover the downstream task [43, 44]. Within the augmentation policy scenario bKLD(small) outperformed all other task selectors consistently on all meta metrics (see Fig. 5), demonstrating the strength of its image representation. In the co-training scenario, where the unweighted large bKLD variant performed best, it appears intuitive that focusing primarily on one task’s dominant features does not yield better transfer estimation. Though the bKLD(small) and bKLD(large) variants’ meta metrics were very close to each other in most cases (see Fig. 5, 6, Supplementary Fig. LABEL:fig:app_comparison).
While the knowledge cloud may hide the actual task fingerprints from users by returning only distances and metadata, the central service will always see the full task fingerprints. To preserve privacy, it is necessary to sufficiently obfuscate patient information within the fingerprints to prevent attacks such as membership inference or data reconstruction [45]. bKLD offers three mechanisms to accomplish this: aggregation, decoupling, and quantization. The binning procedure is key here. In this step feature vectors of individual images are broken up (i.e., decoupled or shuffled), and each entry is assigned to separate bins (i.e., quantized or perturbed). Moreover, the final fingerprint consists of the aggregated histograms for all images. Although we do not provide full theoretical evidence, we can point to the literature, where these techniques are common tools in differential privacy [46].
Consideration of computational costs is highly relevant in the context of both democratization of AI and environmental impact. Related to this work, two different perspectives need to be distinguished: The resources for developing and validating the present approach and the resources required for using our method. The “brute force” approach of combining every potential knowledge transfer between two tasks of our task pool was required to assess the meta metrics weightedtau, percentile and (most of the time) regret. But these are not the costs incurred during regular use: Computing task fingerprints (i.e., feature extraction and binning) took us only about 2 minutes for all tasks together on a standard workstation with 12 CPU cores and a single RTX 3090. A regular participant has to compute only a single task fingerprint once, which stored as a numpy array, makes about 0.4 MB (100 bins) to 4.1 MB (1000 bins). For every conducted experiment during model development, the outcome and hyperparameters for the corresponding fingerprint can be submitted in an automated fashion, without any noticeable overhead. The most compute intensive part remains the task distance computation in the cloud during a query, where all previously submitted fingerprints need to be considered. Fortunately, computing B (number of bins) times the Kullback-Leibler Divergence is a “cheap” operation and the computations for each existing fingerprint can be parallelized. In our implementation, computing distances for all combinations of task pairs within the task pool of size 71 took about 4 seconds for the 1000 bins bKLD variant (including data transfer to GPU). For an actual query not all internal combinations must be computed, hence our measurement roughly corresponds to a usual query with a knowledge cloud of a couple thousand distinct task submissions (except for the GPU data transfer). This computation time scales linearly with the size of the knowledge cloud. So on standard consumer hardware and without further potentially existing optimizations, our implementation should already scale to 100,000s of distinct fingerprints.
Closely related to task fingerprinting is the field of Automated Machine Learning (AutoML), which aims to fully automate the configuration of ML algorithms, specifically steps such as data preprocessing, feature engineering, and hyperparameter optimization [47]. It shares the goal of AI democratization by resolving tricky steps in ML and empowers non-experts to generate AI systems. However, AutoML often raises computational expenses, especially in approaches such as Neural Architecture Search (NAS) [14, 48], and, in contrast to our approach, lacks a collaborative mechanism. Similarly, frameworks such as nnU-net [49] have shown promise in offering self-configuring solutions for biomedical image segmentation tasks. Although training adjustment in nnU-net is way faster and substantially less resource-intensive than traditional AutoML, it is limited to the static rules designed by the authors (i.e., no new knowledge is gained automatically over time), focuses on a specific task type (only segmentation) and excludes crucial aspects of the training pipeline (e.g., architecture template and data augmentation are fixed). Federated learning [50], on the other hand, offers a path for collaborative learning while preserving data privacy, but imposes additional burdens related to synchronization, trust establishment, and preparation [51]. Task fingerprinting combines the efficiency of nnU-net with the flexibility of AutoML, while leveraging collaborative efforts such as Federated Learning. Similarly to the latter, a general limitation of task fingerprinting is that the system is vulnerable to malicious actors submitting false data in a public setting. This requires the development of preventive measures, such as detection algorithms [52] or trust scores [53]. In addition, the success of the overall concept relies on the willingness of researchers to share knowledge within the community. Perhaps most importantly and in contrast to all three mentioned approaches, however, our method has the potential to facilitate Lifelong Learning [54] by continuously and automatically growing knowledge, as more data and tasks are incorporated over time. This capability could lead to long-term scalability of AI which would reduce energy consumption and carbon footprint of systems [55].
It is important to consider our findings in the context of previous research on task distance measures. In line with our preliminary work [31] the FED task selector ranks high for the model architecture and pretraining scenarios and some meta metrics. However, it falls short on the other meta metrics and knowledge transfer scenarios. Prior studies have also highlighted the inconsistency of transferability estimations in general [37] and specifically in the medical domain [56]. This aligns well with our observations regarding the low stability scores of existing distance measures across target tasks and the transfer scenario and emphasizes the importance of finding a flexible approach that can adapt to multiple transfer scenarios. In fact, Standley et al. [57], also observed that task similarity measures might be suited for one kind of knowledge transfer, e.g., pretraining, but inappropriate for another, e.g., co-training. From this perspective, the consistency with bKLD is remarkable. bKLD also excels as a task selector for 3D segmentation tasks outside of our original experimental design. Our aggregation over multiple metrics and meta metrics was intended to counteract the sensitivity of evaluations to specific metrics [58] and increase the assessment robustness as well as allow for the consideration of multiple knowledge transfer scenarios. We observed that weightedtau allows for the most stable separation of task selectors, as opposed to regret, percentile, and improvement (Fig. 5 a , Supplementary Fig. LABEL:fig:app_comparison). However, it is hard to interpret and circumvents a resource limitation-based cutoff.
The emergence of foundation models [8] has amplified the importance of AI democratization, as the necessary expertise, computational resources, and data volumes have grown substantially. These models could bridge the gap between different modalities in medical imaging through their large-scale pretraining [59]. However, two fundamental problems accompany them: first, selecting the most suitable one from the diverse landscape of available ones , and second, devising an effective adaptation strategy for specific downstream tasks. Although integrating multiple modalities into fingerprints is a topic for future research, we note that task fingerprinting is a promising approach to addressing these challenges. Notably, the computational overhead of task fingerprinting – both in terms of archiving post-experimental knowledge and querying prospective knowledge – is minimal, especially compared to the actual model training process. Therefore, task fingerprinting has the potential to accelerate the selection and adaptation of foundation models.